# Closing the stellar labels gap: Stellar label independent evidence for [/M] information in Gaia BP/RP spectra

### v1

@@ Abstract
Data-driven models for stellar spectra which depend on stellar labels suffer from label systematics which decrease model performance: the “stellar labels gap”. To close the stellar labels gap, we present a stellar label independent model for Gaia BP/RP (XP) spectra. We develop a novel implementation of a variational auto-encoder; a scatter VAE, which learns to generate an XP spectrum and intrinsic scatter without relying on stellar labels. We demonstrate that our model achieves competitive XP spectra reconstructions in comparison to stellar label dependent models. We find that our model learns stellar properties directly from the data itself. We then apply our model to XP/APOGEE giant stars to study the [/M] information in Gaia XP. We provide strong evidence that the XP spectra contain meaningful [/M] information by demonstrating that our model learns the -bimodality without relying on stellar label correlations, for stars with 5000 K. We publicly release our trained model, codebase and data. Importantly, our stellar label independent model can be implemented for any/all XP spectra because our model performance scales with training object density, not training label density.
@@ 1 Introduction
The advent of increasingly large-scale spectroscopic surveys, such as APOGEE (Majewski et al., 2017), LAMOST (Yan et al., 2022), GALAH (Buder et al., 2021), and most recently Gaia BP/RP (XP; Gaia Collaboration, 2022), has motivated astronomers to develop data-driven models to cope with the massive influx of data (e.g. Ness et al., 2015; Ting et al., 2019; O’Briain et al., 2021; Leung & Bovy, 2019; Zhang et al., 2023; Leung & Bovy, 2023; Li et al., 2023). These data-driven techniques often seek to discern stellar properties from stellar spectra and/or generate stellar spectra from stellar properties.
@@
Data-driven models have begun to be favoured over their physics-driven predecessors due to a disconnect between theory and practice known as the “synthetic gap.” The synthetic gap is a combination of theoretical systematics and instrumental effects which in tandem produce practically unresolvable differences between synthetic and observed stellar spectra, including but not limited to: one-dimensional modeling, assumption of hydrostatic equilibrium and local thermal equilibrium, and telluric lines. For a more comprehensive overview, see the Introduction of O’Briain et al. (2021).
However, it is seldom emphasized that any data-driven model which relies on stellar labels, a term which typically refers to effective temperature surface gravity metallicity [M/H] (and occasionally -abundance /M]), is implicitly physics-driven. Indeed, any stellar labels utilized during training of such a data-driven model must have at some point been estimated from theoretical stellar spectra. Furthermore, this ab initio physics-driven estimation is often times several generations removed from the data-driven model being implemented, as it is becoming increasingly more common to ‘train machine learning on machine learning.’111In this context, training a machine learning model on stellar labels which were themselves obtained from a machine learning model.
We therefore introduce the concept of an additional gap: the “stellar labels gap,” ecompassing stellar label systematics which negatively impact performance of data-driven models which rely on stellar labels (stellar label dependent models). The stellar labels gap includes:
@@
Stellar multiplicity222While in principle stellar label models can account for binarity, in practice this is rarely implemented. (binaries, triples, etc.)
@@
To that end, this work presents a stellar label independent model for Gaia XP spectra. We develop a novel implementation of a variational auto-encoder; a scatter VAE, which learns to generate an XP spectrum while simultaneously estimating intrinsic scatter for individual XP spectra. We demonstrate certain advantages of our stellar label independent approach by comparing our model performance to stellar label dependent models. Subsequently, we interpret the behaviour of model by contrasting stellar label space to our latent space. We then apply our model to the high- and low- sequences to provide stellar label independent evidence that the Gaia XP spectra contain meaningful [/M] information.
The subsequent Cections of this paper are organized as follows: Section 2 briefly reviews the data used in this work, namely Gaia XP spectra and APOGEE stellar labels. Section 3 presents our stellar label independent model architecture as well as our model training procedure. Section 4 compares our trained model performance to stellar label dependent models and interprets what our model has learned about the XP spectra. We then use our model, in Section 5 to conclusively demonstrate that the Gaia XP spectra contain meaningful [/M] information, without the well known issue of -abundance correlations with stellar labels, for stars with 5000 K. Finally, in Section 6 we conclude by discussing the implications of our work in the context of stellar label in/dependent modeling of stellar spectra, its limitations, and promising future applications.
@@ 2.1 Gaia low-resolution BP/RP spectra
The Gaia low-resolution XP spectra in Gaia Data Release 3 (GDR3, Gaia Collaboration, 2022) is comprised of 220+ million flux-calibrated, low-resolution spectra (De Angeli et al., 2022; Montegriffo et al., 2022). These spectra combine measurements from the Blue Photometer (BP) and Red Photometer (RP) Gaia instruments which span 330-680 and 640-1050 nm, respectively.
We use the XP coefficient spectra in this work. XP spectra deviate from traditional spectra because they are typically reported in Hermite polynomial space. Specifically, the BP and RP spectra are transformed from discrete wavelength-space into continuous coefficient-space: 110 coefficients which weight a set of Hermite polynomials (55 blue and 55 red coefficients). One can transform coefficients into fluxes, and vice-versa, ideally without loss of information. Previous machine learning models have been successfully implemented when using data in both XP wavelength space and coefficient space. For instance, Leung & Bovy (2023) train their large astronomy model in XP coefficient space, whereas Zhang et al. (2023) train their deep stellar label model in XP wavelength-space.
@@ 2.2 Catalogs
To construct our train/test datasets, we perform a cross-match betweeen the GDR3 XP spectra and the astroNN333https://github.com/henrysky/astroNN value-added stellar label catalog for APOGEE Data Release 17 (DR17, Leung & Bovy, 2019; Abdurro’uf, 2022). In principle, our model can be trained on the entire XP dataset, since our model does not require stellar labels. However, a primary focus of this work is to compare our stellar label independent model to stellar label dependent models, as well as compare generative spaces: stellar label space versus latent space. As such, we restrict ourselves to the cross-matched XP/APOGEE dataset. For the purposes of this work, we construct three different catalogs from the XP/APOGEE cross-match: the full catalog, the good labels catalog and the pristine labels catalog:
@@
We compare the Kiel diagrams of the three catalogs in Figure 2. The full catalog appears to have poorly estimated labels at both ends of the main sequence, and the metallicity gradient along the giant branch is somewhat obfuscated. Conversely, the good labels catalog has both a ‘cleaner’ main sequence and ‘sharper’ giant branch metallicity gradient. Finally, the pristine catalog is exclusively composed of giant stars with very accurate stellar labels.
@@ 2.3 Preprocessing
Before training our model on the full catalog, we perform the following pre-processing on the XP coefficient spectra. We first adopt the now common practice of concatenating the 55 BP and 55 RP coefficients into a single vector of length 110. Second, we normalize each individual spectrum by its corresponding Gaia band mean flux. Finally, we standard normalize the XP spectra coefficient by coefficient to zero mean and unit variance. Recently, Zhang et al. (2023) implemented median/quantile normalization for XP preprocessing to reduce the negative impact of outliers during training. However, one of the strengths of our model is the ability to incorporate outliers through our star-by-star scatter estimation, described in Section 3.2. As such, we opt for standard normalization.
@@ 3.1 Variational auto-encoder review
An AE, which need not be variational, can be thought of as a non-linear generalization of Principal Component Analysis. An AE begins with an encoder which compresses input data, in our case a stellar spectrum, to a low-dimensional latent representation. A decoder then attempts to reconstruct the stellar spectrum from the latent representation. Ideally, this will allow the latent representation to learn key features which are shared across a set of stellar spectra. The variational nature of a VAE is added to an AE by upgrading the latent space from a collection of discrete points to a latent distribution . The most popular VAE methodology is that of Kingma & Welling (2013), who encode input data onto an independent multivariate Gaussian distribution. The latent space can therefore be entirely characterized by a latent mean vector and variance vector with latent space dimension
@@ 3.2 Scatter variational auto-encoder
We present the high-level architecture of our model in Figure 3. Our sVAE differs from a traditional VAE since, in addition to a decoder (green) which estimates stellar spectra, we introduce a second ‘decoder’ which estimates intrinsic scatter on a star-by-star basis (purple). Importantly, both the XP reconstruction and XP scatter estimate are generated from the same latent space (red). After discarding the encoder (orange) post-training, new XP coefficients (scatter) can be generated given an arbitrary latent space vector.
The input of our encoder is a pre-processed XP spectrum. The XP coefficients are fed through five intermediate layers, composed of 90, 70, 50, 30 and 10 neurons, respectively. All intermediate layers are activated by the gaussian error linear unit (GELU). The latent parameters are then given by a linear transformation of the final intermediate layer. In this work, we have fixed the latent space to 6 dimensions, meaning the encoder produces 12 outputs (6 means and variances). Note that as is typical with VAEs, the reparametrization trick randomly samples (standard normal distribution) to transform the latent mean and variance vectors into a latent space vector via Our decoder is the mirror image of our encoder, and takes as input a vector drawn from the latent distribution. We then reconstruct an XP spectrum, by feeding the latent vector through five intermediate layers analogous to the encoder, except in reverse order. Finally, a reconstruction of the 110 XP coefficients is produced with a linear transform. The scatter estimator has the same architecture as the decoder, except we enforce positivity with a final Sigmoid activation. Importantly, weights and biases of the scatter estimator are entirely disconnected from the decoder. The scatter estimator does not produce an estimate of intrinsic scatter in the traditional sense; variance of the entire XP dataset assuming zero measurement error. Rather, the intrinsic scatter is ‘intrinsic’ to an individual star, because the scatter estimator produces different outputs on a star-by-star, or latent vector-by-latent vector basis. As such, it is more accurate to think of the scatter estimate as an error term which includes traditional intrinsic scatter, systematics, outliers, etc.
@@ 3.3 Training
After data pre-processing (see Section 2.3), we train our model on 90% full catalog, and allocate the remaining 10% for test data. We adopt a similar training approach to Leung & Bovy (2023), and use the AdamW optimizer (Kingma & Ba, 2014; Loshchilov & Hutter, 2017) with Cosine Annealing with Warm Restarts (Loshchilov & Hutter, 2016) as our learning-rate scheduler. Cosine Annealing is a non-monotic learning rate schedule with cosine-shaped cycles. We select an initial cycle learning rate of and a final learning rate of We train our sVAE for 5000 epochs, with 10 cycles of 500 epochs. As in Leung & Bovy (2023), we find that a batch size of 1024 is ideal for balancing training speed and model performance.
@@
where () are the latent means (variances) and we are summing over the 6 latent space dimensions. In the above form, is not particularly intuitive. For the purposes of latent space structure loss, it can be thought as a distance between probability distributions. if two distributions are identical and otherwise. Hence, assuming a multivariate normal latent distribution, the KL divergence is a measure of the Gaussianity of the sVAE latent space.
@@ 3.4 Data and code
Our sVAE model, implemented in PyTorch, can be trained on a single NVIDIA RTX4070 Ti GPU. Our model can project XP spectra into the latent space, and simulate XP spectra from the latent space in a few seconds. Our trained stellar label independent model and codebase are publicly available at https://github.com/AlexLaroche7/xp_vae for others to reproduce our results, build upon our existing model and apply our model to XP spectra beyond the XP/APOGEE cross-match. Furthermore, all data associated with this work is available at https://zenodo.org/records/10951393.
@@ 4.1 Model performance
Large astronomy model (Leung & Bovy, 2023, LB23): A transformer-based model which transfers the methods of Large Language Models (LLMs) to astronomical data: a ‘large astronomy model’ (LAM). LB23 can perceive/predict any combination of XP data or stellar labels. As such, we compare with LB23 by implementing their model in both a stellar independent and dependent way. Specifically:
@@
Stellar label dependent: We project the XP coefficients into the embedding space, estimate stellar labels from the embedding space, re-project the labels into the embedding space, and finally reconstruct the coefficients. The stellar labels for LB23 are: , and [M/H] and and colours from 2MASS photometry (Skrutskie et al., 2006). Near-infrared photometry is included to provide a proxy for extinction by breaking the temperature-extinction degeneracy.
@@
Deep stellar label model (Zhang et al., 2023, ZGR23): A data-driven model which estimates stellar parameters. ZGR23 can also generate XP spectra from , [M/H], extinction and distance. Note that because ZGR23 produce their own stellar parameter estimates, we use the ZGR23 stellar parameters catalog to generate XP spectra with their model. Additionally, ZGR23 advise that their results are only reliable for stars which pass their reliability cut, which we enforce by requiring quality_flags 8 (see Section 4 for further discussion on the implications of their reliability cut).
We compare our model reconstruction errors over test data (10% of the full catalog) to errors from both stellar label independent models. LB23 and ZGR23 are similar in the sense that they both generate XP spectra from labels, but dissimilar in their output: LB23 predicts XP spectra in coefficient space, whereas ZGR23 predicts XP spectra in wavelength space. In order to compare to ZGR23, we convert our model XP coefficient predictions to wavelength space with GaiaXPy.444https://gaia-dpci.github.io/GaiaXPy-website/, version 2.1.0: 10.5281/zenodo.6558515
@@
where is integrating the square of over coefficients. Note that because we are summing over the first 32 BP and RP coefficients due to the length 64 context window of LB23.
@@
where is integrating the square of over wavelengths. Note that because ZGR23 predict the flux at 61 XP wavelengths.
@@
In Figure 5, we compare the relative coefficient error distributions for our model to LB23 across the first five BP and RP coefficients (Eq. (6)). Our relative error distributions are Gaussian with negligible bias, relative to the stellar label dependent LB23 model. Conversely, the stellar label independent LB23 model evidently outperforms our model over most coefficients. Additionally, in Figure 6, we compare the ratio of coefficient reconstruction errors over the full catalog for LB23 relative to our trained sVAE (Eq. (7)), for both the stellar label independent and dependent implementations. We find that, for the lowest order XP coefficients, our model outperforms the stellar label dependent LB23 model by 1-2 orders of magnitude. Furthermore, the relative performance increase for our model decays to unity for higher order coefficients, which are the most noise dominated. Conversely, we observe that the stellar label independent LB23 model outperforms our sVAE by an order of magnitude for all but the lowest order coefficients.
@@
We conclude that our sVAE outperforms the stellar label dependent model, but not the stellar label independent model, of LB23. This is expected, given the length 64 context window of LB23. In other words, the stellar label independent LB23 model is effectively learning a one-to-one mapping between coefficients, without data compression. In contrast, our model compresses the XP spectrum into 6 latent variables, and as such suffers from a significant amount of XP information loss relative to the LB23 stellar independent model.
@@
Analogously, in Figure 7, we compare the relative wavelength error distributions for our model to ZGR23 across 10 wavelengths uniformly distributed throughout the XP spectra wavelength range (Eq. (8)). Then, in Figure 8, we compare the ratio of wavelength reconstruction errors for ZGR23 relative to our model (Eq. (9)) Here, we only compare errors for stars in the full catalog which satisfy the ZGR23 reliability cut, which decreases the test sample from 50,232 stars down to 37,302 stars. Our stellar label independent model produces less bias from 600-800 nm (in the red half of the XP spectra). Both the wavelength reconstruction errors and relative wavelength error distributions suggest that our model better reconstructs XP spectra for most wavelengths beyond the blue end, relative to ZGR23. By inspecting the model wavelength error ratios in Figure 8, we observe that our model outperforms ZGR23 in flux reconstruction for most of the wavelengths from 550-900 nm by approximately 20-40%. However, for wavelengths at the extremities ZGR23 achieves equivalent or better performance than our sVAE, concentrated towards the blue end ( nm). Lastly, there are certain anomalous wavelengths which do not appear to follow the general error ratio trend as a function of wavelength.

### v2

@@ Abstract
Data-driven models for stellar spectra which depend on stellar labels suffer from label systematics which decrease model performance: the āstellar labels gapā. To close the stellar labels gap, we present a stellar label independent model for Gaia BP/RP spectra. We develop a novel implementation of a variational auto-encoder, which learns to generate an XP spectrum and accompanying āscatterā without relying on stellar labels. We demonstrate that our model achieves competitive XP spectra reconstructions in comparison to stellar label dependent models. We find that our model learns stellar properties directly from the data itself. We then apply our model to XP/APOGEE giant stars to study the [/M] information in Gaia XP. We provide strong evidence that the XP spectra contain meaningful [/M] information by demonstrating that our model learns the -bimodality, without relying on stellar label correlations for stars with 5000 K, while also being sensitive to the anomalous abundances of Gaia-Enceladus stars. We publicly release our trained model, codebase and data. Importantly, our stellar label independent model can be implemented for any/all XP spectra because our model performance scales with training object density, not training label density.
@@ 1 Introduction
The advent of increasingly large-scale spectroscopic surveys, such as APOGEE (Majewski etĀ al., 2017), LAMOST (Yan etĀ al., 2022), GALAH (Buder etĀ al., 2021), and most recently Gaia BP/RP (XP; Gaia Collaboration, 2022), has motivated astronomers to develop data-driven models to cope with the massive influx of data (e.g. Ness etĀ al., 2015; Ting etĀ al., 2019; OāBriain etĀ al., 2021; Leung & Bovy, 2019; Zhang etĀ al., 2023; Leung & Bovy, 2023; Li etĀ al., 2023). These data-driven techniques often seek to discern stellar properties from stellar spectra and/or generate stellar spectra from stellar properties.
@@
Data-driven models have begun to be favoured over their physics-driven predecessors due to a disconnect between theory and practice known as the āsynthetic gap.ā The synthetic gap is a combination of theoretical systematics and instrumental effects which in tandem produce discrepancies between synthetic and observed stellar spectra, including but not limited to: one-dimensional modeling, assumption of hydrostatic equilibrium and local thermal equilibrium, and telluric lines. Although significant progress has been achieved in stellar atmosphere modeling, e.g. three-dimensional non-local thermal equilibrium models (Bergemann, 2014), these simplified modeling assumptions are still frequently adopted when determining stellar properties of large-scale surveys. For a more comprehensive overview, see the Introduction of OāBriain etĀ al. (2021).
However, it is seldom emphasized that any data-driven model which relies on stellar labels, a term which typically refers to effective temperature surface gravity metallicity [M/H] (and occasionally -abundance /M]), is implicitly physics-driven. Indeed, stellar labels utilized during training of such a data-driven model are typically estimated from theoretical stellar spectra. This is not strictly the case: e.g. and for Gaia FGK benchmark stars were determined with angular diameter measurements and bolometric fluxes (Heiter etĀ al., 2015). However, such samples contain nowhere near the requisite number of stars for training, although useful for validation (34 Gaia FGK benchmark stars). Furthermore, this ab initio physics-driven estimation is often times several generations removed from the data-driven model being implemented, as it is becoming increasingly more common to ātrain machine learning on machine learning.ā In this context: training a machine learning model on stellar labels which were themselves obtained from a machine learning model. It has been demonstrated that training large language models on synthetic data produces a decrease in output diversity which worsens with each successive āML on MLā iteration (Guo etĀ al., 2023). A similar concern for stellar properties is therefore warranted.
We therefore introduce the concept of an additional gap: the āstellar labels gap,ā ecompassing stellar label systematics which negatively impact performance of data-driven models which rely on stellar labels (stellar label dependent models). The stellar labels gap includes:
@@
Stellar multiplicity111While in principle stellar label models can account for binarity, in practice this is rarely implemented. (binaries, triples, etc.)
@@
To that end, this work presents a stellar label independent model for Gaia XP spectra: an unsupervised learning model which applies data compression. We develop a novel implementation of a variational auto-encoder (VAE); a scatter VAE, which learns to generate an XP spectrum while simultaneously estimating intrinsic scatter for individual XP spectra. We demonstrate certain advantages of our stellar label independent approach by comparing our model performance to stellar label dependent models. Subsequently, we interpret the behaviour of model by contrasting stellar label space to our latent space. We then apply our model to the high- and low- sequences to provide stellar label independent evidence that the Gaia XP spectra contain meaningful [/M] information.
The subsequent Sections of this paper are organized as follows: Section 2 briefly reviews the data used in this work, namely Gaia XP spectra and APOGEE stellar labels. Section 3 presents our stellar label independent model architecture as well as our model training procedure. Section 4 compares our trained model performance to stellar label dependent models and interprets what our model has learned about the XP spectra. We then use our model, in Section 5 to conclusively demonstrate that the Gaia XP spectra contain meaningful [/M] information, without the well known issue of -abundance correlations with stellar labels, for stars with 5000 K. Finally, in Section 6 we conclude by discussing the implications of our work in the context of stellar label in/dependent modeling of stellar spectra, its limitations, and promising future applications.
@@ 2.1 Gaia low-resolution BP/RP spectra
The Gaia low-resolution XP spectra in Gaia Data Release 3 (GDR3, Gaia Collaboration, 2022) is comprised of 220+ million flux-calibrated, low-resolution spectra (De Angeli etĀ al., 2022; Montegriffo etĀ al., 2022). These spectra combine measurements from the Blue Photometer (BP) and Red Photometer (RP) Gaia instruments which span 330-680 and 640-1050 nm, respectively.
We use the XP coefficient spectra in this work. XP spectra deviate from traditional spectra because they are typically reported in Hermite polynomial space. Specifically, the BP and RP spectra are transformed from discrete wavelength-space into continuous coefficient-space: 110 coefficients which weight a set of Hermite polynomials (55 blue and 55 red coefficients). One can transform coefficients into fluxes, and vice-versa, ideally without loss of information. Previous machine learning models have been successfully implemented when using data in both XP wavelength space and coefficient space. For instance, Leung & Bovy (2023) train their large astronomy model in XP coefficient space, whereas Zhang etĀ al. (2023) train their deep stellar label model in XP wavelength-space.
@@ 2.2 Catalogs
To construct our train/test datasets, we perform a cross-match betweeen the GDR3 XP spectra and the astroNN222https://github.com/henrysky/astroNN value-added stellar label catalog for APOGEE Data Release 17 (DR17, Leung & Bovy, 2019; Abdurroāuf, 2022). In principle, our model can be trained on the entire XP dataset, since our model does not require stellar labels. However, a primary focus of this work is to compare our stellar label independent model to stellar label dependent models, as well as compare generative spaces: stellar label space versus latent space. As such, we restrict ourselves to the cross-matched XP/APOGEE dataset. For the purposes of this work, we construct three different catalogs from the XP/APOGEE cross-match: the full catalog, the good labels catalog and the pristine labels catalog:
ā¢
@@
We compare the Kiel diagrams of the three catalogs in Figure 2. The full catalog appears to have poorly estimated labels at both ends of the main sequence, where surface gravities are systematically overestimated, and the metallicity gradient along the giant branch is somewhat obfuscated. Conversely, the good labels catalog has both a ācleanerā main sequence and āsharperā giant branch metallicity gradient. The pristine catalog is then exclusively composed of giant stars with very accurate stellar labels. Finally, apart from the stellar label cuts defining the pristine catalog removing all main-sequence stars, we do not observe major differences between the apparent magnitude and Gaia color distributions across our catalogs, beyond the color and magnitude cuts we impose.
@@ 2.3 Preprocessing
Before training our model on the full catalog, we perform the following pre-processing on the XP coefficient spectra. We first adopt the now common practice of concatenating the 55 BP and 55 RP coefficients into a single vector of length 110. Second, we normalize each individual spectrum by its corresponding Gaia band mean flux. Finally, we standard normalize the XP spectra coefficient by coefficient to zero mean and unit variance. Recently, Zhang etĀ al. (2023) implemented median/quantile normalization for XP preprocessing to reduce the negative impact of outliers during training. However, one of the strengths of our model is the ability to incorporate outliers through our star-by-star scatter estimation, described in Section 3.2. As such, we opt for standard normalization.
@@ 3.1 Variational auto-encoder review
An AE, which needs not be variational, can be thought of as a non-linear generalization of Principal Component Analysis. An AE begins with an encoder which compresses input data, in our case a stellar spectrum, to a low-dimensional latent representation. A decoder then attempts to reconstruct the stellar spectrum from the latent representation. Ideally, this will allow the latent representation to learn key features which are shared across a set of stellar spectra. The variational nature of a VAE is added to an AE by upgrading the latent space from a collection of discrete points to a latent distribution . The most popular VAE methodology is that of Kingma & Welling (2013), who encode input data onto an independent multivariate Gaussian distribution. The latent space can therefore be entirely characterized by a latent mean vector and variance vector with latent space dimension
@@ 3.2 Scatter variational auto-encoder
We present the high-level architecture of our model in Figure 3. Our sVAE differs from a traditional VAE since, in addition to a decoder (green) which estimates stellar spectra, we introduce a second ādecoderā which estimates intrinsic scatter on a star-by-star basis (purple). Importantly, both the XP reconstruction and XP scatter estimate are generated from the same latent space (red). After discarding the encoder (orange) post-training, new XP coefficients (and scatter) can be generated given an arbitrary latent space vector.
The input of our encoder is a pre-processed XP spectrum. The XP coefficients are fed through five intermediate layers, composed of 90, 70, 50, 30 and 10 neurons, respectively. All intermediate layers are activated by the gaussian error linear unit (GELU). The latent parameters are then given by a linear transformation of the final intermediate layer. In this work, we have fixed the latent space to 6 dimensions, meaning the encoder produces 12 outputs (6 means and variances). This choice was informed by initial experimentation with latent dimensions ranging from 1 to 20 dimensions. We found that a 6 dimensional latent space produced an optimal balance between interpretability and reconstruction error. Note that as is typical with VAEs, the reparametrization trick randomly samples (standard normal distribution) to transform the latent mean and variance vectors into a latent space vector via Our decoder is the mirror image of our encoder, and takes as input a vector drawn from the latent distribution. We then reconstruct an XP spectrum, by feeding the latent vector through five intermediate layers analogous to the encoder, except in reverse order. Finally, a reconstruction of the 110 XP coefficients is produced with a linear transform. The scatter estimator has the same architecture as the decoder, except we enforce positivity with a final Sigmoid activation. Importantly, weights and biases of the scatter estimator are entirely disconnected from the decoder.
The interpretation of what the scatter estimator represents is by no means straightforward. The scatter estimator does not produce an estimate of intrinsic scatter in the traditional sense: variance of the entire XP dataset assuming zero measurement error. Rather, the scatter could be both āintrinsicā to an individual star: stellar variability arising from stellar spots, pulsations, etc. or simply be anomalous relative to the training set, but also āextrinsicā: arising from either information loss through our data compression procedure or Gaia systematics, poor observations or underestimation of uncertainties. As such, it is more accurate to think of the scatter estimate as an error term which includes traditional intrinsic scatter, systematics, outliers, etc. Empirically, we noticed a marked increase in performance when including the scatter estimator in our model. Therefore, despite not being able to pinpoint precisely which of the above aspects the scatter estimate mitigates, we can assert that our sVAE architecture does improve spectral modeling.
@@ 3.3 Training
After data pre-processing (see Section 2.3), we randomly split the full catalog into 90% for training, and allocate the remaining 10% for test data. We adopt a similar training approach to Leung & Bovy (2023), and use the AdamW optimizer (Kingma & Ba, 2014; Loshchilov & Hutter, 2017) with Cosine Annealing with Warm Restarts (Loshchilov & Hutter, 2016) as our learning-rate scheduler. Cosine Annealing is a non-monotic learning rate scheduler with cosine-shaped cycles. We select an initial cycle learning rate of and a final learning rate of We train our sVAE for 5000 epochs, with 10 cycles of 500 epochs. As in Leung & Bovy (2023), we find that a batch size of 1024 is ideal for balancing training speed and model performance.
@@
where () are the latent means (variances) and we are summing over the 6 latent space dimensions. In the above form, is not particularly intuitive. For the purposes of latent space structure loss, it can be thought as a distance between probability distributions. if two distributions are identical and otherwise. Hence, assuming a multivariate normal latent distribution, the KL divergence is a measure of the Gaussianity of the sVAE latent space. Note that no variables directly appear in Eq. (5) due to marginalizing out from the joint probability distribution
Finally, we emphasize that Eq. (1) does not account for the full Gaia XP coefficient covariance matrix. Our decision to neglect covariances during training is based on computational feasibility. Specifically, the inclusion of the full covariance matrix would have lead to an increased computational cost due to both loss function evaluation and the scatter output necessarily being augmented to two dimensions (to match the covariance matrix) if we took this approach.
@@ 3.4 Data and code
Our sVAE model, implemented in PyTorch, can be trained on a single NVIDIA RTX4070 Ti GPU in 6 hours, with our current XP/APOGEE cross-match of training objects333Linearly extrapolating this training time to the full Gaia DR3 XP dataset of 220 million spectra, we roughly estimate that training on all currently available XP spectra would take 120 GPU days (on the same GPU, with the same batch size, architecture, etc.). To give some perspective, a foundation language model such as LLaMA (Touvron etĀ al., 2023) takes 21 days 2048 GPUs = GPU days to train.. Our model can project XP spectra into the latent space, and simulate XP spectra from the latent space in a few seconds. Our trained stellar label independent model and codebase are publicly available at https://github.com/AlexLaroche7/xp_vae for others to reproduce our results, build upon our existing model and apply our model to XP spectra beyond the XP/APOGEE cross-match. Furthermore, all data associated with this work is available at https://zenodo.org/records/10951393.
@@ 4.1 Model performance
Large astronomy model (Leung & Bovy, 2023, LB23): A transformer-based model which transfers the methods of Large Language Models (LLMs) to astronomical data: a ālarge astronomy modelā (LAM). LB23 can perceive/predict any combination of XP data or stellar labels. As such, we compare with LB23 by implementing their model in both a stellar independent and dependent way. Specifically:
@@
Stellar label dependent: We project the XP coefficients into the embedding space, estimate stellar labels from the embedding space, re-project the labels into the embedding space, and finally reconstruct the coefficients. The stellar labels for LB23 are: , and [M/H] and and colours from 2MASS photometry (Skrutskie etĀ al., 2006). Near-infrared photometry is included to provide a proxy for extinction by breaking the temperature-extinction degeneracy.
@@
Deep stellar label model (Zhang etĀ al., 2023, ZGR23): A data-driven model which estimates stellar parameters. ZGR23 can also generate XP spectra from , [M/H], extinction and distance. Note that because ZGR23 produce their own stellar parameter estimates, we use the ZGR23 stellar parameters catalog to generate XP spectra with their model. Additionally, ZGR23 advise that their results are only reliable for stars which pass their reliability cut, which we enforce by requiring quality_flags 8 (see Section 4 for further discussion on the implications of their reliability cut).
We compare our model reconstruction errors over test data (10% of the full catalog) to errors from both stellar label independent models. LB23 and ZGR23 are similar in the sense that they both generate XP spectra from labels, but dissimilar in their output: LB23 predicts XP spectra in coefficient space, whereas ZGR23 predicts XP spectra in wavelength space. In order to compare to ZGR23, we convert our model XP coefficient predictions to wavelength space with GaiaXPy.444https://gaia-dpci.github.io/GaiaXPy-website/, version 2.1.0: 10.5281/zenodo.8239995 (Ruz-Mieres & Kostrzewa-Rutkowska, 2023)
@@
where computes the reconstruction error while accounting for the XP covariance matrices. Specifically, and are vectors containing the first 32 BP and RP coefficients due to the length 64 context window of LB23. Similarly, and are the model reconstructions for the first 32 BP and RP coefficients, and () are the 3232 sub-matrices of the full BP (RP) covariance matrices.
@@
where is integrating the square of over wavelengths. Note that because ZGR23 predict the flux at 61 XP wavelengths555GaiaXPy does not currently have the functionality to account for the full XP covariance matrices, and as such we only consider univariate errors in Eq. (9)..
@@
In Figure 5, we compare the relative coefficient error distributions for our model to LB23 across the first five BP and RP coefficients, Eq. (6). Our relative error distributions are Gaussian with negligible bias, relative to the stellar label dependent LB23 model. Conversely, the stellar label independent LB23 model evidently outperforms our model over most coefficients. Additionally, in Figure 6, we compare the ratio of coefficient reconstruction errors over the full catalog for LB23 relative to our trained sVAE, Eq. (7), for both the stellar label independent and dependent implementations. We find that, for the lowest order XP coefficients, our model outperforms the stellar label dependent LB23 model by 1-2 orders of magnitude. Furthermore, the relative performance increase for our model decays to unity for higher order coefficients, which are the most noise dominated. Conversely, we observe that the stellar label independent LB23 model outperforms our sVAE by an order of magnitude for all but the lowest order coefficients.
@@
We conclude that our sVAE outperforms the stellar label dependent model, but not the stellar label independent model, of LB23. This is expected, given the length 64 context window of LB23. In other words, the stellar label independent LB23 model is effectively learning a one-to-one mapping between coefficients, without data compression. In contrast, our model compresses the XP spectrum into 6 latent variables, and as such loses XP information relative to the LB23 stellar independent model.
@@
Analogously, in Figure 7, we compare the relative wavelength error distributions for our model to ZGR23 across 10 wavelengths uniformly distributed throughout the XP spectra wavelength range, Eq. (8). Then, in Figure 8, we compare the ratio of wavelength reconstruction errors for ZGR23 relative to our model,Eq. (9). Here, we only compare errors for stars in the full catalog which satisfy the ZGR23 reliability cut, which decreases the test sample from 50,232 stars down to 37,302 stars. Our stellar label independent model produces less bias from 600-800 nm (in the red half of the XP spectra). Both the wavelength reconstruction errors and relative wavelength error distributions suggest that our model better reconstructs XP spectra for most wavelengths beyond the blue end, relative to ZGR23. By inspecting the model wavelength error ratios in Figure 8, we observe that our model outperforms ZGR23 in flux reconstruction for most of the wavelengths from 550-900 nm by approximately 20-40%. However, for wavelengths at the extremities ZGR23 achieves equivalent or better performance than our sVAE, concentrated towards the blue end ( nm). Here, it is informative to compare the APOGEE Kiel diagram in Figure 2 to the training sample of ZGR23 (Figure 1 in Zhang etĀ al., 2023). We speculate that the larger number of hot stars in the ZGR23 training sample may contribute to our sVAE under-performing at bluer wavelengths. Lastly, there are certain anomalous wavelengths which do not appear to follow the general error ratio trend as a function of wavelength.